OpenStudy (anonymous):

What are to numbers whose difference is 16 and is as small as possible?

OpenStudy (tkhunny):

TWO numbers. A and B, A being the greater. Difference is 16. A - B = 16 Product is A*B = ????? Only guessing at what it wanted. Please provide a better problem statement and your best work.,

OpenStudy (anonymous):

Let x = one number x - 16 = send number (whose difference is 16) You want the product x(x-16) = x^2 - 16x to be as small as possible. Pretty straightforward...as you can find the minimum point of this parabola (via x = -b/2a, the line of symmetry passing through the minimum of the parabola, etc.).

OpenStudy (tkhunny):

...assuming, of course, that we have any idea at all what the actual problem is.

OpenStudy (anonymous):

The actual problem is to find two numbers whose difference is 16 and whose product is a minimum.

OpenStudy (tkhunny):

Who told you that? Nothing like a magic problem statement to make life interesting.

OpenStudy (anonymous):

Thats how I read the problem, just the author of the problem spelled two as to...but it was clear to me what he meant.

OpenStudy (anonymous):

Perhaps I read it wrong.....